### v1

# Krylov Complexity as a Probe for Chaos

Abstract
In this work, we explore in detail, the time evolution of Krylov complexity. We demonstrate, through analytical computations, that in finite many-body systems, while ramp and plateau are two generic features of Krylov complexity, the manner in which complexity saturates reveals the chaotic nature of the system. In particular, we show that the dynamics towards saturation precisely distinguish between chaotic and integrable systems. For chaotic models, the saturation value of complexity reaches its infinite time average at a finite saturation time. In this case, depending on the initial state, it may also exhibit a peak before saturation. In contrast, in integrable models, complexity approaches the infinite time average value from below at a much longer timescale. We confirm this distinction using numerical results for specific spin models.
I Introduction
While we have a solid understanding of classical chaos, quantum chaos remains a more elusive concept. Classical chaotic behaviour is characterized by the sensitivity of trajectories in phase space to initial conditions, where two trajectories initially close to each other diverge exponentially fast — a phenomenon quantified by the Lyapunov exponent. In contrast, quantum chaos is a more ambiguous concept and is rather challenging to understand. This is primarily due to the absence of a clear understanding of quantum phase space.
In an isolated quantum system defined by its Hamiltonian, along with specific boundary or initial conditions, it is natural to expect that the key indicators of its chaotic nature are encoded in the structure of its eigenstates and eigenvalues. Indeed, it is well known that the eigenvalues of chaotic Hamiltonians display distinct statistical features. More precisely, the level spacing of a maximally chaotic quantum system follows the Wigner-Dyson distribution, while that of integrable systems follows the Poisson distribution Bohigas et al. (1984).
To investigate the nature of quantum chaos, several quantities have been introduced, including out-of-time- order correlators (OTOCs) Shenker and Stanford (2014a, b). Semiclassically, OTOCs exhibit exponential growth due to the butterfly effect, characterized by the Lyapunov exponent, which is conjectured to be bounded Maldacena et al. (2016). This bound is saturated in certain strongly interacting models with holographic descriptions, such as the Sachdev-Ye-Kitaev (SYK) model Kitaev ; Sachdev and Ye (1993); Maldacena and Stanford (2016). However, it is important to note that the exponential growth of OTOCs is not a universal characteristic of all chaotic systems Fine et al. (2014); Xu and Swingle (2019); Hashimoto et al. (2020).
The present article aims to explore the chaotic nature of many-body quantum systems using Krylov complexity [^1] Throughout this paper, we will examine the Krylov complexity of states Balasubramanian et al. (2022); Alishahiha and Banerjee (2023). , as defined in the context of the Lanczos algorithm. While this method has traditionally been employed for numerical computations Viswanath and Müller (1994), in recent years there has been increased interest in applying the Krylov method to the study of quantum chaos Parker et al. (2019).
The authors of Parker et al. (2019) proposed a universal operator growth hypothesis that connects the asymptotic behaviour of Lanczos coefficients to the dynamics of the system in question. Specifically, for chaotic many-body quantum systems in dimensions greater than one and lacking symmetry, the Lanczos coefficients are expected to exhibit asymptotic linear growth. However, while the literature supports this linear growth in chaotic systems, it may not be a universal indicator of chaos; non-chaotic models can also display asymptotic linear growth Dymarsky and Smolkin (2021); Bhattacharjee et al. (2022a); Avdoshkin et al. (2024); Camargo et al. (2023); Vasli et al. (2024).
The Krylov complexity saturates in many-body systems when the Lanczos coefficients vanish at certain steps. It is therefore natural to examine this saturation value to determine whether it can distinguish between chaotic and integrable models. Actually, in an explicit Ising model, it has been demonstrated that the saturation value of the Krylov complexity increases as one transitions from an integrable model to a chaotic one Rabinovici et al. (2022a, b). However, it is important to note that the saturation value of Krylov complexity also depends on the specific state for which it is computed. In fact, the saturation value for chaotic systems may or may not exceed that of integrable models Rabinovici et al. (2022a, b); Scialchi et al. (2024); Ballar Trigueros and Lin (2022); Español and Wisniacki (2023). We will return to this point later.
Using the evolution of maximally entangled states in the Krylov basis, it was shown in Erdmenger et al. (2023) that the linear growth and saturation of Krylov complexity are generic features of many-body systems, regardless of whether they are chaotic or integrable. However, in chaotic systems, the Krylov complexity exhibits a peak before reaching its saturation value at late times Balasubramanian et al. (2022).
The presence of this peak before saturation has recently been observed numerically in various models Huh et al. (2024); Camargo et al. (2024); Nandy et al. (2024); Bhattacharjee and Nandy (2024); Balasubramanian et al. (2024); Baggioli et al. (2024). It is important to note that the appearance of a peak prior to saturation is not solely determined by the system’s dynamics; it also depends on the initial state. For instance, it occurs for the initial state considered in Erdmenger et al. (2023). Nonetheless, it appears that the thermofield double state at infinite temperature provides a maximal universal bound on this peak Camargo et al. (2024).
In this paper, we aim to further explore how Krylov complexity can probe the chaotic nature of systems. To achieve this, we leverage the fact that Krylov complexity is essentially the expectation value of the number operator in the Krylov basis. Consequently, studying the time dependence of complexity reduces to analyzing the time evolution of this expectation value—an object we are more familiar with in the context of quantum thermalization (see e.g. Alishahiha (2023)).
Inspired by the procedure that leads to the eigenstate thermalization hypothesis (ETH) Srednicki (1994); Deutsch (1991), we find that the growth and subsequent saturation of complexity at a time exponentially large in relation to the system size is a generic feature of Krylov complexity, regardless of whether the dynamics is chaotic or integrable. The specific aspect of complexity that is sensitive to the chaotic nature of the system is how it approaches its saturation value at late times. In chaotic systems, the saturation value of complexity reaches the infinite time average of the complexity at the saturation time, whereas in integrable models, it approaches this value from below over a relatively longer time.
It is worth noting that while the aforementioned proposal outlines the generic features of complexity for chaotic and integrable models, the detailed behaviour of complexity also depends on the initial state. In the chaotic case, for a maximally entangled state, the Krylov complexity exhibits a clear peak before reaching its saturation value. However, for generic initial states, this peak may or may not exist, or it may not be pronounced. Additionally, there are small fluctuations around the saturation value, the shape and nature of which also depend on the initial state.
The rest of the paper is organized as follows. In section II, we briefly review the Lanczos algorithm and present a closed analytic form for the expansion coefficient of Krylov basis in terms of the energy eigenstates, for a general closed many body system. This makes the perfect background on which in section III we study several generic features of Krylov complexity leading to the main claim of this paper. In section IV we examine our proposal using numerical results from an explicit many-body system. The last section is devoted to discussions.
II Krylov basis vs energy eigenbasis
To set the stage, let us consider a closed quantum system with a time-independent local non-degenerate Hamiltonian $H$ whose eigenstates and eigenvalues are denoted by $|E_{\alpha}\rangle$, and $E_{\alpha}$, respectively. Here $\alpha=1,\cdots,{\cal D}$ with ${\cal D}$ being dimension of full Hilbert space. Starting with an initial state, $|\psi_{0}\rangle$ at $t=0$, its time evolution in the Schrödinger picture is given by
$|\psi(t)\rangle=e^{-iHt}|\psi_{0}\rangle\,.$ (1)
At $t=0$, the initial state can be expanded in terms of the energy eigenbasis as
$|\psi_{0}\rangle=\sum_{\alpha=1}^{\cal D}c_{\alpha}|E_{\alpha}\rangle\,.$ (2)
Starting with the same initial state $|\psi_{0}\rangle$ in the same quantum system with time-independent Hamiltonian $H$, the Krylov basis, $\{|n\rangle,n=0,1,2,\cdots,{\cal D}_{\psi}-1\}$ can be constructed as follows. The first element of the basis is identified with the initial state $|0\rangle=|\psi_{0}\rangle$ which we assume to be normalized. The other elements are then constructed, recursively, following the algorithm Viswanath and Müller (2008)
$(H-a_{n})|n\rangle=b_{n+1}|n+1\rangle+b_{n}|n-1\rangle\,,$ (3)
where $a_{n}=\langle n|H|n\rangle$, and
$b^{2}_{n+1}=\langle n|H^{2}|n\rangle-a^{2}_{n}-b_{n}^{2}\,\;\;\;\;\;{\rm with}\;b_{0}=0.$ (4)
Note that, physically, $a_{n}$ is the average energy associated with the state $|n\rangle$. Moreover, denoting the variance of energy for this state as $\Delta a_{n}$ with
$(\Delta a_{n})^{2}=\langle n|H^{2}|n\rangle-\langle n|H|n\rangle^{2}\,,$ (5)
one gets
$b^{2}_{n+1}=\sum_{j=0}^{n}(-1)^{j}(\Delta a_{n-j})^{2}\,.$ (6)
This recursive procedure stops whenever $b_{n}$ vanishes which occurs for $n={\cal D}_{\psi}\leq{\cal D}$ which defines the dimension of Krylov space. Note that this algorithm produces an orthonormal and ordered basis together with coefficients $a_{n}$ and $b_{n}$ known as Lanczos coefficients Lanczos (1950).
Once we have constructed the Krylov basis, the evolved state at any time can be expanded in this basis
$|\psi(t)\rangle=\sum_{n=0}^{{\cal D}_{\psi}-1}\phi_{n}(t)\,|n\rangle\,,\;\;\;\;{\rm with}\;\;\sum_{n=0}^{{\cal D}_{\psi}-1}|\phi_{n}(t)|^{2}=1\,,$ (7)
where the dynamics of the wave function $\phi_{n}(t)$ is governed by the Schrödinger equation
$-(i\partial_{t}+a_{n})\phi_{n}(t)=b_{n}\phi_{n-1}(t)+b_{n+1}\phi_{n+1}(t)\,,$ (8)
which should be solved using the initial condition $\phi_{n}(0)=\delta_{n0}$.
An advantage of working with the Krylov basis lies in the fact that under a unitary time evolution, the trajectory of a given initial state does not necessarily spread into the entire Hilbert space. Instead, it remains confined within a subset known as the Krylov space. The latter typically has a smaller dimension compared to the full Hilbert space of the system. Thus, following the argument above, the Krylov space suffices for studying the time evolution of the system. In particular, if the system has conserved charges, working in this basis we are automatically confined in a subsystem that preserves the symmetry of the initial state.
To elaborate on the last point, let us expand an element of the Krylov space in terms of the energy eigenbasis
$|n\rangle=\sum_{\alpha=1}^{{\cal D}}f_{n\alpha}|E_{\alpha}\rangle\,.$ (9)
Note that since ${\cal D}_{\psi}\leq{\cal D}$, the expansion coefficient, $f_{n\alpha}$, is not necessarily invertible and therefore, in general, energy eigenstates cannot be expanded in terms of Krylov basis. Using the orthogonality condition of the Krylov basis one gets
$\sum_{\alpha=1}^{{\cal D}}f^{*}_{n\alpha}f_{m\alpha}=\delta_{nm}\,.$ (10)
However, it is important to note that, on general grounds, the condition ${\cal D}_{\psi}\leq{\cal D}$ arises from a symmetry in the model and this can be understood as follows. Let us assume our model has a symmetry generated by the charge $Q$. Given this symmetry, we have $[Q,H]=0$. This means that starting with an initial state of a specific symmetry charge, the entire Krylov basis is constrained to maintain that same charge.
This can be argued by noting that the energy eigenstates can actually be organized into different blocks according to their symmetry charges. Therefore, selecting a particular initial state with a definite charge, its spreading through the Krylov basis nevertheless retains that same charge. The dimension of the Krylov space coincides with that of the corresponding symmetry block Herviou et al. (2021). As an immediate consequence of this fact, the summation in the equation (9) terminates at ${\cal D}_{\psi}$
$|n\rangle=\sum_{j=1}^{{\cal D}_{\psi}}f_{nj}|E_{j}\rangle\,,\;\;\;\;\;\;\;{\rm for}\;\;n\leq{\cal D}_{\psi}-1.$ (11)
In this case, $f_{nj}$ can be inverted within a fixed charge sector and accordingly, in this symmetry block, the energy eigenvectors can also expanded in terms of the Krylov basis. Note that, within this sector, one still has the orthogonality relation,
$\sum_{j=1}^{{\cal D}_{\psi}}f^{*}_{nj}f_{mj}=\delta_{nm}\,.$ (12)
Following this justification, we will consider an initial state with a definite symmetry charge, which restricts the summation over energy eigenstates in (2) to the fixed charge subsystem with dimension ${\cal D}_{\psi}$, the dimension of Krylov subspace, leading to an expansion
$|\psi_{0}\rangle=|0\rangle=\sum_{j=1}^{{\cal D}_{\psi}}c_{j}|E_{j}\rangle.$ (13)
Plugging (11) in (3) yields [^2] In our notation, $\{n,m\}=0,\cdots{{\cal D}_{\psi}-1}$ while $\{i,j\}=1,\cdots{{\cal D}_{\psi}}$. The Greek indices $\{\alpha,\beta\}$ are reserved for the full Hilbert space, namely, $\{\alpha,\beta\}=1,\cdots{\cal D}$.
$(E_{j}-a_{n})f_{nj}=b_{n+1}f_{n+1\,j}+b_{n}f_{n-1\,j}\,,$ (14)
for $n\leq{\cal D}_{\psi}-1$, that can be solved to find $f_{nj}$ in terms of $f_{0j}=c_{j}$ as
$f_{nj}=\frac{g_{n}(E_{j})}{b_{n}!}c_{j}\,.$ (15)
Here we use the notation that $b_{n}!=b_{1}\cdots b_{n}$ and $b_{0}!=1$, and $g_{n}(E_{j})={\rm det}(E_{j}-H^{(n)})$, $H^{(n)}$ being an $n\times n$ matrix whose elements are given by
$H^{(n)}_{pq}=\langle p|H|q\rangle=a_{q}\delta_{pq}+b_{q+1}\delta_{pq+1}+b_{q}\delta_{pq-1}\,,$ (16)
for $p,q=0,\cdots,n-1$. Substituting (15) into (12) and (14), one obtains an orthogonality relation for the function $g_{n}(E_{i})$ and a recursion relation, respectively, of the forms [^3] Equation (18) can be immediately recognized as the recursion relation of the continuant of a tridiagonal matrix. This is consistent with the fact that Hamiltonian has a tridiagonal form in Krylov basis Balasubramanian et al. (2023). We also observe that this recurrence relation actually resembles the three-term recurrence relation for Krawtchouk polynomials Miki et al. (2022). This relationship can, in turn, be utilized to derive an exact solution for $a_{n}$ and $b_{n}$.
$\sum_{i=1}^{{\cal D}_{\psi}}|c_{i}|^{2}g_{n}(E_{i})g_{m}(E_{i})=\left(b_{n}!\right)^{2}\,\delta_{nm}\,,$ (17)
$(E_{i}-a_{n})g_{n}(E_{i})=g_{n+1}(E_{i})+b_{n}^{2}g_{n-1}(E_{i})\,.$ (18)
Note that, by definition, $g_{-1}(E_{j})=0,g_{0}(E_{j})=1$. Furthermore, using the definition of wave function in terms of the Krylov basis (7) one can find the function $\phi_{n}$ in terms of $g_{n}(E_{i})$ as
$\phi_{n}(t)=\frac{1}{b_{n}!}\sum_{j=1}^{{\cal D}_{\psi}}|c_{j}|^{2}g_{n}(E_{j})\,e^{-iE_{j}t}\,.$ (19)
It is straightforward to see that by plugging this expression into equation (8) one arrives at (18).
III Krylov complexity probes chaos
Endowed with the general relations derived in the previous section, in the present section, we establish the main claim of our paper, namely, we explore the key role of Krylov complexity in detecting the chaotic behaviour of a given system at a late time.
Krylov complexity is defined as the expectation value of the number operator in the Krylov basis Parker et al. (2019)
${\cal C}(t)=\langle\psi(t)|{\cal N}|\psi(t)\rangle\,,$ (20)
where the number operator in the Krylov basis is defined as
${\cal N}=\sum_{n=0}^{{\cal D}_{\psi}-1}\,n|n\rangle\langle n|\,.$ (21)
To investigate the time dependence of complexity, we must therefore examine the time evolution of the expectation value (21). At the end of this section, we will be able to demonstrate that this expectation value, at late times, will be an important marker to detect the chaotic nature of any given system.
To proceed, let us consider, once again, the initial state (13). With this, the complexity (20) reads
${\cal C}(t)={\rm Tr}(\rho_{\rm DE}{\cal N})+\sum_{j\neq k}^{{\cal D}_{\psi}}e^{i\omega_{jk}t}c^{*}_{j}c_{k}{\cal N}_{jk}\,,$ (22)
where $\omega_{jk}=E_{j}-E_{k}$, ${\cal N}_{jk}=\langle E_{j}|{\cal N}|E_{k}\rangle$ and the diagonal density matrix $\rho_{\rm DE}$ given by
$\rho_{\rm DE}=\sum_{j=1}^{{\cal{D}}_{\psi}}|c_{j}|^{2}|E_{j}\rangle\langle E_{j}|\,.$ (23)
It is worth emphasizing once more, starting from the initial state (13), we only focus on a fixed charge sector of the full Hilbert space spanned by the Krylov basis. Accordingly, the diagonal density matrix $\rho_{\rm DE}$ defined in (23) is confined to that particular symmetry-resolved block only.
The infinite time average of complexity is then given by
[…]
Knowing the explicit form of the number operator allows us to extract generic features of Krylov complexity. For this, it is useful to write an explicit form of the matrix elements of the number operator. Indeed, from equation (11) one finds
${\cal N}_{jk}=\sum_{n=0}^{{\cal D}_{\psi}-1}nf_{nj}f^{*}_{nk}\,.$ (25)
Thereafter, using (15), one arrives at
${\cal N}_{jk}=c_{j}c^{*}_{k}\sum_{n=0}^{{\cal D}_{\psi}-1}\frac{ng_{n}(E_{j})g_{n}(E_{k})}{\left(b_{n}!\right)^{2}}\,.$ (26)
Furthermore, by definition, the complexity is zero at $t=0$, so that, from (22) we have
${\cal C}(0)=\sum_{j,k=1}^{{\cal D}_{\psi}}c^{*}_{j}c_{k}{\cal N}_{jk}=0\,.$ (27)
This leads to
$\overline{{\cal C}}=\sum_{j}^{{\cal D}_{\psi}}|c_{j}|^{2}{\cal N}_{jj}=-\sum_{j\neq k}^{{\cal D}_{\psi}}c^{*}_{j}c_{k}{\cal N}_{jk}>0\,.$ (28)
For finite many-body systems where the Lanczos coefficients vanish at the step $n={\cal D}_{\psi}$, the Krylov complexity saturates [^4] It is worth noting that for models where $b_{n}$ vanishes at $n={\cal D}_{\psi}$, one may have a symmetry so that $b_{n}=b_{{\cal D}_{\psi}-n}$. In this case, the vanishing of the Lanczos coefficients does not result in complexity saturation, instead, we get a periodic behaviour, which could be due to an underlying $SU(2)$ symmetry Caputa et al. (2022). Indeed, the exact solution of equation (18) based on Krawtchouk polynomials belongs to this category. Of course, in what follows, we will consider cases without such a symmetry, so that the complexity saturates at late times. . Therefore, at late times, we have
${\cal C}(t)\approx{\cal C}_{\rm S}+{\rm small\;fluctuations}\,,$ (29)
where ${\cal C}_{\rm S}$ is the saturation value of complexity. To achieve such a behaviour, the time-dependent phase in equation (22) must sum to zero at late times, leading to a constant value equal to the late-time saturation value of the complexity. This constant should be contrasted with the infinite time average. Actually, from (22) one finds
$\overline{|{\cal C}(t)-\overline{{\cal C}}|^{2}}=\sum_{j\neq k}^{{\cal D}_{\psi}}|c_{j}|^{2}|c_{k}|^{2}{\cal N}^{2}_{jk}\,,$ (30)
which demonstrates that $\overline{{\cal C}}\approx{\cal C}_{\rm S}$ iff the right-hand side of (30) is sufficiently small. In the context of quantum thermalization, this is always the case for operators that obey the ETH ansatz, where the off-diagonal matrix elements are exponentially small. This also ensures a rapid thermalization of any typical ETH observable right after the system has relaxed. On the contrary, it is important to note that the number operator is not an ETH operator. Consequently, it is not immediately clear whether the saturation value corresponds to the infinite time average of complexity.
From (28) it is clear that the off-diagonal matrix elements of the number operator are of the same order as the diagonal ones. Therefore, in generic many-body systems with exponentially dense energy levels, the expression from (22) suggests that if complexity saturates, the onset of saturation should occur at a time that is exponentially large in terms of the size of the system. This in turn makes the non-ETH-like behaviour of the number operator manifest.
In conclusion, we observe that from equation (22), for generic many-body systems, the Krylov complexity begins at zero value and continues to grow until it reaches an exponentially large time, at which point it saturates at a value which is of the order of the infinite time average of Krylov complexity. Notably, this behaviour is a general characteristic of Krylov complexity in all many-body systems, regardless of whether they are integrable or chaotic.
From equation (22) and following the observation that complexity saturation begins at exponentially large times, we can actually estimate the saturation value of complexity. Interestingly enough, unlike the previous general observation, the saturation value of complexity does actually depend on whether the system is chaotic or integrable, as we explore in what follows.
To proceed, we use the explicit form of the matrix elements of the number operator to rewrite the complexity (22) as follows [^5] To arrive at this result, we have used the fact that the expression for complexity obtained by plugging (26) into (22), is manifestly symmetric under the exchange of $j$ and $k$.
${\cal C}(t)=-2\sum_{n=0}^{{\cal D}_{\psi}-1}\frac{n}{\left(b_{n}!\right)^{2}}\sum_{j\neq k}^{{\cal D}_{\psi}}\sin^{2}(\frac{\omega_{jk}}{2}t)|c_{j}|^{2}|c_{k}|^{2}g_{n}(E_{j})g_{n}(E_{k})\,.$ (31)
From this expression, it is evident that the complexity is zero at $t=0$ as expected. Furthermore, it also manifests the expected quadratic growth at early times. Finally, it saturates at late times. To determine the saturation value, in what follows we examine the behaviour of complexity at large times.
Since we are interested in the complexity at the large time limit, it is useful to arrange the energy eigenvalues in increasing order, $E_{1}<E_{2}<\cdots<E_{{\cal D}_{\psi}}$, and reformulate the summation in equation (31) as follows
${\cal C}(t)=-2\sum_{n=0}^{{\cal D}_{\psi}-1}\frac{n}{\left(b_{n}!\right)^{2}}\sum_{\ell=1}^{{\cal D}_{\psi}-1}\sum_{j=1}^{{\cal D}_{\psi}-\ell}\sin^{2}(\frac{s_{j+\ell}}{2}t)$ (32)
×$\hskip 88.2037pt\times\,|c_{j+\ell}|^{2}|c_{j}|^{2}g_{n}(E_{j+\ell})g_{n}(E_{j})\,,$
where $s_{j+\ell}=E_{j+\ell}-E_{j}$. It is clear that the large-time behaviour of complexity depends on the distribution of $s_{j+\ell}$, which, in turn, relies on the distribution of energy eigenvalues. This is where the information about the nature of the system, whether it is chaotic or integrable, enters into the computations.
Since the eigenvalues are arranged in an increasing order, the smallest values of $s_{j+\ell}$, namely the ones with $\ell=1$ oscillate slower than those with higher $\ell$ in the late time limit. The contributions coming from these slow oscillating modes towards the late-time behaviour of complexity are thus determined by the level spacing of the system. It is well known that for integrable systems, the level spacing follows a Poisson distribution, while for chaotic systems, it conforms to the Wigner surmise [^6] In general for chaotic systems we have Wigner-Dyson distribution given by $P(s)=\mu_{0}s^{\delta}e^{-\mu_{1}s^{2}}$, for $\delta=1,2,3$ and ${\cal O}(1)$ constants $\mu_{0}$ and $\mu_{1}$ to normalize the distributions.
$P_{\rm In}(s)=e^{-s},\;\;\;\;\;\;P_{\rm Ch}(s)=\frac{\pi}{2}\,s\,e^{-\frac{\pi}{4}s^{2}}\,.$ (33)
In chaotic systems, the distribution of $s_{j+1}$ exhibits [^7] It is worth noting, however, $s_{j+l}$ appearing in (32) has the dimension of energy while the level spacing $s$ appearing in (33) is dimensionless. Of course, it is known that to find the universal level spacing an appropriate normalization is required Evnin and Piensuk (2019). a peak around a value which is of order one, and this behaviour extends to other $\ell$ as well. The $\sin^{2}$ factor oscillates between zero and one, and for large $t$, its average value approaches $\frac{1}{2}$. Thus, we can estimate the complexity ${\cal C}(t)$ at late times as
${\cal C}(t)\approx-\sum_{n=0}^{{\cal D}_{\psi}-1}\frac{n}{\left(b_{n}!\right)^{2}}\sum_{j\neq k}^{{\cal D}_{\psi}}|c_{j}|^{2}|c_{k}|^{2}g_{n}(E_{j})g_{n}(E_{k})\,.$ (34)
This leads to
${\cal C}(t)\approx\overline{{\cal C}}+{\rm fluctuations}\,,$ (35)
at late times. In order to identify the late time saturation value of complexity with the infinite time average, as in (35), we have used (28). The nature of these subleading fluctuations, however, depends on the specific state for which the complexity is calculated. It is worth emphasizing that the information about the late time saturation of complexity is encoded in the Lanczos coefficients appearing in the definition of $g_{n}(E_{j})$.
On the other hand, for integrable models, the Poisson distribution suggests that the peak of level spacing for $s_{j+1}$ is centred around zero. Nevertheless, for large values of $\ell$ this peak shifts towards an ${\cal O}(1)$ value. Consequently, we need to decompose the complexity into two parts: the first part involves fast oscillations from $s_{j+\ell}$ for $\ell>1$, and the second part, corresponding to $\ell=1$. The latter represents a slowly oscillating term.
For large $t$, we can approximate $\sin^{2}$ by $\frac{1}{2}$ in the first part, leading to a constant contribution. However, since the first term does not consider $\ell=1$, the resultant constant is smaller than the infinite time average of the complexity. Denoting this value as ${\cal C}_{0}<\overline{{\cal C}}$, we find
${\cal C}(t)\approx{\cal C}_{0}-2\sum_{n=0}^{{\cal D}_{\psi}-1}\frac{n}{\left(b_{n}!\right)^{2}}\sum_{j=1}^{{\cal D}_{\psi}-1}\sin^{2}(\frac{s_{j+1}}{2}t)$ (36)
×$\hskip 88.2037pt\times\,|c_{j+1}|^{2}|c_{j}|^{2}g_{n}(E_{j+1})g_{n}(E_{j})\,.$
For sufficiently large $t$, we can still estimate the $\sin^{2}$ of the second term as $\frac{1}{2}$, which results in complexity saturation approaching the infinite time average. However, for intermediate (yet still large) $t$, the second part oscillates, indicating that for integrable models, complexity saturation remains below the infinite time average and approaches it from below as time progresses.
Therefore, the saturation value of the Krylov complexity and the way it saturates in finite many-body systems could serve as a measure to determine whether the model is integrable or chaotic. To be precise, we observe that at the saturation time
${\cal C}_{S}\approx{\rm Tr}(\rho_{\rm DE}{\cal N}),\;\;\;\;\;\;\;\;\;{\rm for\;\;chaotic\;\;systems}\,,$ (37)
${\cal C}_{S}<{\rm Tr}(\rho_{\rm DE}{\cal N}),\;\;\;\;\;\;\;\;\;{\rm for\;\;integrable\;\;systems}\,.$
It is also insightful to compute the time derivative of the complexity around the saturation time, which yields the dominant contribution of the form
$\dot{{\cal C}}(t)\approx-\sum_{n=0}^{{\cal D}_{\psi}-1}\frac{n}{\left(b_{n}!\right)^{2}}\sum_{j=1}^{{\cal D}_{\psi}-1}\sin(s_{j+1}t)\,s_{j+1}$ (38)
×$\hskip 88.2037pt\times\,|c_{j+1}|^{2}|c_{j}|^{2}g_{n}(E_{j+1})g_{n}(E_{j})\,.$
For integrable models, this expression approaches zero due to the $s_{\alpha}$ factor. In contrast, for chaotic models, due to the level spacing distribution and for certain initial states this could in general result in a finite value, which may be almost zero at a particular time $t_{m}$ identified with the saturation time. Moreover, it is also possible to show that the second derivative of this term is negative at that point, indicating that if such an extremum exists, it must be a maximum. Consequently, in this case, the complexity exhibits a peak before saturating at the value determined by the infinite time average of the complexity. Of course, it is important to note that, in general, depending on the initial state, the peak may not be present at all or even if it is present, it may not be pronounced enough.
Thus, a distinguishing signature of whether a model is chaotic or integrable can be signalled through how the Krylov complexity approaches its infinite time average at late times. For integrable models, the complexity approaches the average from below, whereas for chaotic models, it approaches the saturation value at a finite time. Furthermore, depending on the initial state, it may also exhibit a peak before saturation.
We note, however, that for certain initial states of a chaotic system, the complexity does exhibit a clear peck before saturation. To elaborate on this point further, we note that for any initial state $|\psi_{0}\rangle$ in a typical many-body system, one can define the canonical density matrix as follows
$\rho_{\rm th}=\frac{e^{-\beta H}}{Z(\beta)},\;\;\;\;Z(\beta)={\rm Tr}\left(e^{-\beta H}\right)\,,$ (39)
where the inverse temperature $\beta$ is determined by the equation ${\rm Tr}(\rho_{\rm th}{H})=\langle\psi_{0}|H|\psi_{0}\rangle=E_{0}$. It is also useful to define the thermal Krlyov complexity as
${\cal C}_{\rm th}={\rm Tr}(\rho_{\rm th}{\cal N})\,,$ (40)
which could be the saturation value of Krylov complexity for the initial state $|\psi_{0}\rangle$ if the number operator were an ETH operator.
For chaotic systems, one would expect that expectation values of typical operators saturate to their thermal values shortly after relaxation, though as we have argued before, we would expect complexity to saturate at exponentially long times. Therefore, the saturation value of Krylov complexity should exceed the thermal Krylov complexity, which is attained at the thermalization time. This is a simple consequence of the fact that the thermalization time is significantly shorter than the time required for the complexity to saturate.
With these definitions in mind, let us now consider the following initial state
$|\psi_{0}\rangle=\frac{1}{\sqrt{Z(\beta)}}\sum_{j=1}^{{\cal D}_{\psi}}e^{-\frac{1}{2}\beta E_{j}}|E_{j}\rangle\,.$ (41)
For this state, interestingly, the infinite time average of complexity equals the thermal Krylov complexity
${\cal C}(t)={\cal C}_{\rm th}+\frac{1}{Z(\beta)}\sum_{j\neq k}^{{\cal D}_{\psi}}e^{-\frac{1}{2}\beta(E_{j}+E_{k})}e^{i(E_{j}-E_{k})t}{\cal N}_{jk}\,.$ (42)
As mentioned above, for chaotic systems, the complexity exceeds the thermal Krylov complexity around the thermalization time and subsequently saturates to its infinite time average at longer times. For the initial state (41), the time average coincides with the thermal Krylov complexity, leading to the observation that the complexity exhibits a peak around the saturation time, approaching the infinite time average from above. In contrast, for integrable models, the complexity nearly reaches the thermal Krylov complexity and approaches its infinite time average from below at late times.
Thus, in a typical chaotic system the complexity can indeed exhibit a clear peak before it saturates to the value given by its infinite time average, for certain initial states, the state given in (41) being one such designated initial state.
IV Explicit examples
In this section, we put our proposal to rigorous numerical tests using a particular many-body system, namely, the spin-$\frac{1}{2}$ Ising model described by the Hamiltonian
$H=-J\sum_{i=1}^{N-1}\sigma^{z}_{i}\sigma^{z}_{i+1}-\sum_{i=1}^{N}(g\sigma^{x}_{i}+h\sigma^{z}_{i})\,.$ (43)
Here $\sigma^{z,y,z}$ are Pauli matrices and $J,g$ and $h$ are constants which define the model. By rescaling one can set $J=1$. Then the nature of the model, namely whether it is chaotic or integrable, is solely determined by the value of the constants $g$ and $h$. In particular, for $gh\neq 0$ the model is non-integrable. In order to perform our numerical computations we will set $h=0.5,\,g=-1.05$ for chaotic case Bañuls et al. (2011), and $h=0,\,g=-1.05$ to extract behaviour in the integrable regime.
After fixing the above-mentioned values of the constants $J,g$ and $h$, we numerically compute the Krylov complexity in this model for several initial states. We specifically consider three different initial states corresponding to configurations where all spins are aligned along the $x,y$ and $z$ directions, denoted as $|X+\rangle,|Y+\rangle,|Z+\rangle$, respectively.
In the chaotic limit of this Ising model (43), these states have been considered in Bañuls et al. (2011) to explore the nature of thermalization, in particular, whether the thermalization is strong or weak [^8] By making use of certain measures the nature of thermalization for different spin models has been studied in Sun et al. (2021a); Lin and Motrunich (2017); Chen et al. (2021); Bhattacharjee et al. (2022b); dos Prazeres and de Oliveira (2024); Alishahiha and Vasli (2024).. In the strong thermalization, the expectation value of a typical operator relaxes to the thermal value very fast, while for weak thermalization it strongly oscillates around the thermal value, though its time average attains the thermal value. It has been shown that although the initial state $|Y+\rangle$ exhibits strong thermalization, the initial states $|Z+\rangle$ and $|X+\rangle$ do not (see also Sun et al. (2021b)).
It is known that the Hamiltonian (43) has a parity symmetry which is essentially the reflection symmetry about the centre of the chain. Therefore, the energy eigenstates can be organized into two blocks according to their parity, whether it is positive or negative. For $N=11$ which is the number of spin we will consider in our numerical computations, the dimension of subspace with the positive party is 1056 and that with negative parity is 992.
It is easy to check that the aforementioned initial states have positive parity which in turn indicates that the obtained Krylov subspace in chaotic case should be a subspace with positive parity. One can compute the dimension of Krylov space associated with these states to obtain ${\cal D}_{\psi}=1056$. It is, however, important to note that for the integrable cases, besides the parity, due to extra possible conserved charges, the dimensions of Krylov spaces are actually smaller than this. In fact, one finds ${\cal D}_{Y,Z}=925$ and ${\cal D}_{X}=463$.
It is straightforward to compute the Lanczos coefficients for these initial states from which one can compute complexity numerically999 Lanczos coefficients for the model under consideration have also been computed inNoh (2021); Español and Wisniacki (2023); Trigueros and Lin (2022); Scialchi et al. (2024); Bhattacharya et al. (2023, 2022) . by using (19). The numerical results for the chaotic ($h=0.5$) and the integrable ($h=0$) cases are depicted in figures 1 and 2, respectively.
Figure 1: Time evolution of complexity for different initial states $|X+\rangle,|Y+\rangle$ and $|Z+\rangle$ for the case where the model is chaotic; $g=-1.05,h=0.5$. The dashed lines represent the infinite time average of complexity. The computation is performed for $N=11$.
Figure 2: Time evolution of complexity for different initial states $|X+\rangle,|Y+\rangle$ and $|Z+\rangle$ for the case where the model is integrable; $g=-1.05,h=0.5$. The dashed lines represent the infinite time average of complexity. The computation is performed for $N=11$.
We observe that the numerical results are perfectly compatible with the analytical argument we presented in the previous section. Indeed, the plots show distinctive behaviour of complexity for chaotic and integrable models. While that the complexity saturates to its saturation value given by the infinite time average of complexity at the saturation time for chaotic, in the integrable case it reaches this value from below at large times.
It is worth noting once again that the actual value of the complexity saturation does not provide a measure to probe whether the model is chaotic or integrable. As evident from the results presented in figures 1 and 2, the infinite time average of complexity for the chaotic system is not significantly different from that of the integrable models101010see Rabinovici et al. (2022a, b); Scialchi et al. (2024); Ballar Trigueros and Lin (2022); Español and Wisniacki (2023) for related discussions.
We have also observed a significant fluctuation around the saturation value, though it does not affect the physical behaviour of the complexity. The size and the nature of fluctuation depend on the initial states. Of course, the fluctuations become milder as we move to higher $N$.
Another interesting observation we have made is that although in the chaotic case and for the initial state $|Y+\rangle$ the complexity exhibits a clear peck before the saturation, it is not the case for other initial states. This is in line with the discussion in the previous section where we argued that for a generic initial state, complexity may or may not exhibit the peak. Even in cases where such a peak exists, it may not be pronounced enough. Of course, having a peak is guaranteed for particular initial states such as the thermofield double state as we have shown in the previous section. The second example we showed numerically is the initial state $|Y+\rangle$ in the chaotic regime of the Ising model (43).
In order to crosscheck our analytic claim regarding the TFD state, we have also computed complexity for the initial state (41) for $\beta=0$. The result is shown in figure (3). From this figure one can immediately observe that the complexity for chaotic case does exhibit a peak before saturation. Furthermore, the behaviour of complexity for the integrable case also confirms the other part of our claim, namely it approaches the infinite time average from below and reaches it at late times.
Figure 3: Time evolution of complexity for initial state given by (41) with $\beta=0$. The dashed lines represent the infinite time average of complexity. One observes that in the chaotic case the complexity exhibits a clear peak before saturation. The computation is performed for $N=11$.
Thus far, our numerical computations have been centred on two specific points within the parameter space of the model (43), specifically $(g=-1.05,h=0)$ and $(g=-1.05,h=0.5)$. These points correspond to integrable and nearly maximally chaotic scenarios, leading to Poisson and Wigner distributions for level spacing, respectively. It would be more interesting to explore additional points and calculate the Krylov complexity, which would allow us to determine whether the system tends towards chaotic or integrable behaviours depending on the parametric closeness of the system to the aforementioned points. In pursuit of this, we have also evaluated the Krylov complexity for the initial state $|Y+\rangle$ at different values for $h$, yielding the results depicted in figure (4). Analysis of these figures reveals that as we progress towards higher values of $h$, the system moves away from chaos and leans towards integrability, as expected. This observation further supports our earlier conclusion in the preceding section that the $\sin$-term and level spacing play crucial roles in determining the behaviour of Krylov complexity at late times.
Figure 4: Time evolution of complexity for initial state $|Y+\rangle$. for $h=1,1.5,2,3$. This shows that as we progress towards higher values of $h$, the system moves away from chaos and leans towards integrability. The computation is performed for $N=11$.
Furthermore, it is instructive to examine the level spacing for different values of $h$ which we present in Figure 4. The corresponding level spacings are shown in Figure 5. The latter clearly demonstrates how the level spacing transitions from a Wigner distribution to a Poisson distribution as $h$ increases.
Figure 5: Level spacing for energy eigenvalues of the model (43) with positive parity for different values of $h$ given by $h=1,1.5,2,3$. This shows that as we progress towards higher values of $h$, the distribution moves away from Wigner and leans towards Poisson. To have a better statistic the computation is performed for $N=13$.
V Discussions
In this work, we have explored whether Krylov complexity can effectively probe the chaotic nature of many-body quantum systems. To address this outstanding question, we have utilized the fact that Krylov complexity is defined as the expectation value of the number operator in the Krylov basis. Inspired by the quantum thermalization process, we have examined the time evolution of complexity, observing that it exhibits some general common features which are not specific to the nature of the particular system in consideration. In particular, we show that at exponentially long times, the Krylov complexity saturates (asymptotically) at a value very close to the infinite time average of complexity, regardless of whether the system is chaotic or integrable.
Nevertheless, we demonstrate that the manner in which complexity saturates does depend, significantly, on the dynamical nature of the systems. For a generic initial state in a typical chaotic system, the saturation value of complexity reaches its infinite time average at the saturation time, whereas in integrable models, it approaches the infinite time average value from below over a long duration.
However, it is important to note that the exact value of complexity at the saturation point and its detailed behaviour at large and intermediate times depend not only on the dynamics but also on the initial state used for computing the Krylov complexity. Specifically, when the initial state is a maximally entangled state, the Krylov complexity, while approaching the infinite time average from above, leads to a peak before ultimately saturating at that average value. The appearance of this peak is not universal or state-independent. In fact, for a generic initial state, the complexity may or may not exhibit a clear peak. Moreover, there are small fluctuations around the saturation value whose shapes also depend on the initial state. This confirms the postulates of Susskind (2016); Brown et al. (2017) for complexity saturation.
Our observation regarding the behaviour of Krylov complexity is founded on the explicit form of the number operator within the Krylov basis. The Krylov complexity can be expressed in the following general form:
$\mathcal{C}(t)=\sum_{j\neq k}^{\mathcal{D}\psi}\sin^{2}\left(\frac{\omega_{jk}}{2}t\right)|c_{j}|^{2}|c_{k}|^{2}A(E_{j},E_{k}),$ (44)
where $A(E_{j},E_{k})$ is a smooth function of its arguments which can be read by comparing this equation with (31). It is evident that the late-time behaviour of complexity is dictated by the distribution of $\omega_{jk}=E_{j}-E_{k}$, a quantity that may provide insights into the chaotic behaviour of the system. Importantly, one must ensure that only those energy eigenvalues that lie within a symmetric block are included. Notably, the Lanczos algorithm used to construct the Krylov basis inherently satisfies this constraint.
It is important to highlight that achieving the desired saturation behaviour requires the function $A(E_{j},E_{k})$ to be smooth and not develop poles, a condition that holds for finite $n$. In infinite-dimensional systems, where $n$ can become very large, we can analyze the late-time behaviour of complexity by adopting the continuum limit. This approach enables us to compute the function $A(E_{j},E_{k})$, which often reveals a double-pole structure, resulting in linear growth of complexity at late times Alishahiha (2023); Alishahiha and Banerjee (2023).
Finally, it is worth noting that for a thermal initial state where $|c_{j}|^{2}=e^{-\beta E_{j}}/Z[\beta]$, the Krylov complexity can be expressed in a form which is reminiscent of the spectral complexity introduced in Iliesiu et al. (2022). Therefore, we would expect that the spectral complexity exhibits the same late-time behaviour as that of the Krylov complexity.
acknowledgements
We would like to thank Rathindra Nath Das, Johanna Erdmenger, René Meyer, Reza Pirmoradian, Mohammad Reza Tanhayi and Hamed Zolfi for useful discussions. This work is based upon research founded by Iran National Science Foundation (INSF) under project No 4023620.